At standard temperature and pressure (0 ∘C and 1.00 atm ), 1.00 mol of an ideal gas occupies a volume of 22.4 L. What volume would the same amount of gas occupy at the same pressure and 80 ∘C ?
Express your answer with the appropriate units.
1
Expert's answer
2019-09-03T02:33:51-0400

From ideal gas equation

We know that


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


Where,"P_1,V_1,T_1," are initial pressure,initial volume and initial temprature respectively

and"P_2,V_2,T_2," are final pressure,final volume and final tempreture respectively

Now,


"V_2=\frac{P_1V_1T_2}{P_2T_1}"


"P_1=101325pascals"

"V_1=0.0224m^3"

"T_1=273.15K"

"P_2=P_1"

"T_2=253.15K"


So, "V_2=\frac{101325\times0.0224\times 253.15}{101325\times 273.15}=0.02076m^3"


Volume occupied at "80\degree C" and same pressure will be "20.76l"